A candy store sells malted milk balls for $2.50 per pound, and chocolate-covered
raisins for $3.75 per pound. How many pounds of each should be used to make a 10
pound mixture that sells for $3.00 per pound?

Answer by venugopalramana(3286)

A candy store sells malted milk balls for $2.50 per pound, and chocolate-covered
raisins for $3.75 per pound. How many pounds of each should be used to make a 10
pound mixture that sells for $3.00 per pound?
LET MILK BALLS USED BE M LB.
RAISINS USED =10-M LB.
COST OF MILK BALLS = M*2.5
COST OF RAISINS = (10-M)*3.75=37.5-3.75M
TOTAL COST =2.5M+37.5-3.75M=10*3=30
3.75M-2.5M=37.5-30=7.5
1.25M=7.5
M=7.5/1.25=6
MILK BALLS USED =6 LB.
RAISINS USED =10-6=4 LB.